# Orbital magnetoelectric coupling of three dimensional Chern insulators

### v2

@@ VI.3 Layer Chern number
Let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. What would happen if a non-disruptive in-plane electric field is applied assuming the interlayer coupling is weak? Since 3D Chern insulator is adiabatically connected to trivial stacking of 2D Chern insulators, the response to in-plane electric field (corresponding to off-diagonal $\alpha_{xz}$ or similar components) should behave similarly as in the 2D case. Specifically, if a uniform electric field $\mathcal{E}_{x}$ is applied along the $x$ axis, a transverse anomalous Hall current flowing along the $y$ direction would be induced due to the non-zero Chern number, with $j_{y}=Ce^{2}/(dh)\mathcal{E}_{x}$. The bulk anomalous Hall current would naturally generate orbital magnetization $\mathbf{M}$ via ×$\nabla\times\mathbf{M}=\mathbf{j}$, giving rise to an off-diagonal OME response, with its response coefficient varying linearly in real space, i.e., $\partial_{x}\alpha_{xz}=-Ce^{2}/hd$. Here $e$ is fundamental charge, $h$ is Planck constant, and $d$ is interlayer distance.

### v3

@@ VII Off-diagonal OME response of Chern insulators
Let us consider a 3D Chern insulator with Chern vector $\textbf{C}=(0,0,C_{z})$, where $C_{z}=C(C\in\mathbb{Z},C\neq 0)$ with interlayer distance $d$. What would happen if a non-disruptive in-plane electric field is applied assuming the interlayer coupling is weak? Since 3D Chern insulator is adiabatically connected to trivial stacking of 2D Chern insulators, the response to in-plane electric field (corresponding to off-diagonal $\alpha_{xz}$ or similar components) should behave similarly as in the 2D case. Specifically, if a uniform electric field $\mathcal{E}_{x}$ is applied along the $x$ axis, a transverse anomalous Hall current flowing along the $y$ direction would be induced due to the non-zero Chern number, with $j_{y}=Ce^{2}/(dh)\mathcal{E}_{x}$. The bulk anomalous Hall current would naturally generate orbital magnetization $\mathbf{M}$ via ×$\nabla\times\mathbf{M}=\mathbf{j}$, giving rise to an off-diagonal OME response, with its response coefficient varying linearly in real space, i.e., $\partial_{x}\alpha_{zx}=-Ce^{2}/hd$. Here $e$ is fundamental charge, $h$ is Planck constant, and $d$ is interlayer distance.
VIII First-principle DFT calculations for $\rm{MnBi}_{2}\rm{Te}_{4}$ class with strain
First-principle density functional theory (DFT) calculations for Mn(Bi/Sb)2Te4 [(Bi/Sb)2Te3]n thin film with $n=0,1,2$ are performed with Vienna Ab-initio Simulation Package (VASP) [54] using the projector-augmented wave (PAW) method and the plane-wave basis with an energy cutoff of 450 eV. The Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional [55] is used in combination with GGA+$U$ method to treat the localized $d$ orbitals of Mn, with $U=3\,\rm{eV}$ set in all calculations. The Brillouin zone is sampled by a $16\times 16\times 1$ $\Gamma$-centered Monkhorst-Pack $\mathbf{k}$-point mesh. Structure optimizations are performed with a force criterion of $0.01$ eV/Å.
The in-plane strain along the $a$ axis is applied by scaling the length of the $a$ axis, and then relaxing the atomic positions with the scaled axis fixed. The out-of-plane compression along the $c$ axis are applied as follows: we first reduce the height of the top layer atoms and then, while fixing both the top and bottom layer atoms, relax the atomic positions between these two layers.
VIII.1 Chern number of $\rm{MnBi}_{2}\rm{Te}_{4}$ class
Once the electronic structure is converged, the Bloch states are projected to the Wannier functions of the Bi-$6p$ (or Sb-$5p$), and Te-$5p$ orbitals to generate the tight-binding Hamiltonian for identifying the Chern number [56, 57]. Searching among many possibilities of Mn(Bi/Sb)2Te4[(Bi/Sb)2Te3]n class, it turns out that the best building block for the construction of 3D Chern insulator slabs is three-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ (see Table S4). Then, we can construct the tight-binding Hamiltonian of 3D Chern insulator slab using three-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ as repetition entity. Suppose the adjacent entity layer is separated by a spacer of thickness of single layer $\rm{MnBi}_{2}\rm{Te}_{4}$. After classifying the Wannier centers into their corresponding sublayer in three-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ entity, we retrieve the interlayer hopping between the two outmost sublayers, served as the Hermitian conjugate of the inter-entity layer coupling.
Table S4: Chern numbers of Mn(Bi/Sb)2Te4[(Bi/Sb)2Te3]n-class thin films under different strain magnitudes. “FM” and “AFM” stand for interlayer ferromagnetic and anti-ferromagnetic configurations, respectively.
$\mathrm{system\,(layers\mid magnetism\mid strain\,axis)}$2.0% -1.5% -1.0% -0.5% 0.0% 0.5% 1.0% 1.5% 2.0%
[…]
VIII.2 Band structures of three-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ with strain
Band structures of three-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ with strain are shown in Fig. S6. The indirect gaps of three-layer $\rm{MnBi}_{2}\rm{Te}_{4}$ (3L MBT) in the presence of in-plane strain -2%, 0% and 2% are 29.7 meV, 46.3 meV and 53.0 meV, respectively. Larger gap results in better quantization of the calculated Chern number given the same $\mathbf{k}$ mesh.
Figure S6: Band structures of 3L MBT in the presence of in-plane strain -2%, 0% and 2% and the zero energy indicates chemical potential.
IX Results using another expression of layer-resolved orbital magnetization
Recently, Seleznev et al. [38] have shown that multiple formulations for the local marker of orbital magnetization exist. Those formulations are not equivalent for systems belonging to certain symmetry classes. The authors have identified one specific form of the marker as the most physically meaningful since it can consistently predict the correct hinge current. To validate our results, we derive the expression of layer-resolved orbital magnetization from the local marker proposed in [38]
$M_{z}=\frac{e}{3\hbar A}{\rm Im\,Tr}\{PxQHQyP-QxPHPyQ\}+\frac{e}{3\hbar A}{\rm Im\,Tr}\{H(QyPxQ-PyQxP)\}+\frac{e}{3\hbar A}{\rm Im\,Tr}\{(QyPxQ-PyQxP)H\},$ (S53)
which is an equal weight linear combination of three possible operator orderings in the trace operation.
The derivation of the definition layer-resolved orbital magnetization from Eq. (S53) follows the same algebra as outlined in Sec. I.B. We then check the validity of our proposed quantization rule using the new definition under exactly the same set-up described in the main text. Remarkably, the quantization rule remains valid under this new definition, as shown in the subsequent Tables S5, S6 and S7, which correspond to Tables I, II  and IV  in the main text, respectively.
Table S5: Layer-resolved $\alpha_{zz}^{(l)}$ in units of $e^{2}/h$ obtained by numerical finite difference using the definition Eq. (S53) proposed in [38]. Same parameters are used as for Table S1 and Table I  in the main text.
Layer Index 1 2 3 4 5 6
AA,5 1.9924 0.9971 -0.0010 -0.9990 -1.9943
AB,5 1.9377 0.9525 -0.0448 -1.0430 -2.0282
AA,6 2.5397 1.5444 0.5464 -0.4519 -1.4502 -2.4448
AB,6 2.4957 1.5086 0.5113 -0.4842 -1.4828 -2.4680
Table S6: Difference between $\alpha_{zz}^{(l)}$ of neighboring layers in units of $e^{2}/h$ obtained by numerical finite difference using the definition Eq. (S53) proposed in [38]. Same parameters are used as for Table S1 and Table II  in the main text.
$\alpha_{zz}^{(l+1)}-\alpha_{zz}^{(l)}$ $l=1$ $l=2$ $l=3$ $l=4$ $l=5$
AA,5 -0.9953 -0.9982 -0.9980 -0.9953
AB,5 -0.9852 -0.9974 -0.9982 -0.9852
AA,6 -0.9953 -0.9980 -0.9983 -0.9983 -0.9946
AB,6 -0.9871 -0.9973 -0.9955 -0.9986 -0.9852
Table S7: Difference between neighboring layer-resolved response in units of $e^{2}/h$ in systems with substrate using the definition Eq. (S53). Same parameters are used as for Table IV  in the main text.
$\alpha_{zz}^{(l+1)}-\alpha_{zz}^{(l)}$ $l=1$ $l=2$ $l=3$ $l=4$
$t_{3}=0,t_{4}=0$0.9972 -0.9972 -0.9972 -0.9972
$t_{3}=1/6,t_{4}=0$0.9972 -0.9972 -0.9972 -0.9972
$t_{3}=0.3,t_{4}=0$0.9972 -0.9972 -0.9972 -0.9972
$t_{3}=1/6,t_{4}=3t_{3}/4$0.9954 -0.9978 -0.9978 -0.9955
$t_{3}=0.3,t_{4}=3t_{3}/4$0.9910 -0.9991 -0.9990 -0.9913